OpenStudy (pphalke):

In how many ways can a president, a vice president, and a treasurer be chosen from a pool of 15 candidates?

jimthompson5910 (jim_thompson5910):

how far did you get with this one?

OpenStudy (pphalke):

i just know 15 but i dont know the rest

jimthompson5910 (jim_thompson5910):

imagine we have 3 slots A,B,C A = president B = vice president C = treasurer we have 15 choices for slot A. How many choices for slot B?

OpenStudy (pphalke):

i think 45

jimthompson5910 (jim_thompson5910):

We start with 15 people. Once the president is selected, s/he cannot serve any other post. So they are taken out of contention. We go from 15 to 15-1 = 14 people So hopefully you agree that we have 14 choices for slot B?

OpenStudy (pphalke):

yes

jimthompson5910 (jim_thompson5910):

how many choices for slot C?

OpenStudy (pphalke):

13?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now multiply out the number of choices for each slot (# of choices for slot A)*(# of choices for slot B)*(# of choices for slot C) = (15)*(14)*(13) = ??

OpenStudy (pphalke):

I get 2730

jimthompson5910 (jim_thompson5910):

me too